Find the maximum or minimum value of the function.
f(t) = 14t2 + 112t + 295

Answer by Alan3354(69443)

Find the maximum or minimum value of the function.
f(t) = 14t2 + 112t + 295
------
The max or min is on the Line of Symmetry
The LOS is t = -b/2a = -112/28
t = -4
--------
f(-4) = 14*16 + 112*(-4) + 295
f(-4) = 71
----------
The minimum is 71
It's a minimum since a is positive.